# Practical Secure Delegated Linear Algebra with Trapdoored Matrices

### v2

@@ 0.A.1.1 Matrix Multiplication
One natural case is that with $m=n=l$ and one online step (i.e. a once-off matrix-matrix multiplication). Here, total client cost is $O\left(n^{2}\left(n^{\epsilon}+\nu\right)\right)$ and total server cost is, apart from one $n\times n$-by-$n\times n$ matrix multiplication, $O(n^{\omega}\delta^{\omega-2})$. In particular, our algorithm is $\left(n^{\epsilon}+\frac{\nu}{n},1+O(\delta^{\omega-2})\right)$-remote-efficient, resulting in a multiplicative $n^{1-\epsilon}$ saving for the client (ignoring the additive term in security parameter) and small overhead for the server.
@@ 0.A.1.2 Online Matrix-Vectors
Another natural case is $m=n,l=1$, and $k$ online steps. Here, setup cost is the same $O\left(n^{2}\left(n^{\epsilon}+\nu\right)\right)$ for the client and $O\left(\delta n^{\omega}\right)$ for the server, which amortize across $n^{\epsilon}$ and $\delta n$ online rounds, respectively. The per-multiplication online cost is $O(n(n^{\epsilon}+\nu))$ for the client and, apart from one $n\times n$ matrix-vector multiplication, $O(\delta n^{2})$ for the server. We therefore obtain (after amortization) the same $\left(n^{\epsilon}+\frac{\nu}{n},1+O(\delta^{\omega-2})\right)$-remote-efficiency guarantee and same small overheads. Note that the amortization occurs in a sublinear number of online steps; this holds because the initial multiplication benefits from fast matrix multiplication, while the online steps do not.
@@ 0.A.2 Per-Operation Analysis
Let $\mathfrak{a}(x)$ be the cost of adding/subtracting two $x$-element tensors; $\tilde{\mathfrak{d}}(x)$ be the cost of taking the dot product of two vectors, one of which is $x$-sparse; and $\mathfrak{m}(a,b,c)$ be the cost of multiplying two matrices, one of which is $a\times b$ and the other of which is $b\times c$. All operations are relative to our working ring.
We analyze the performance componentwise and with precise constants, as this algorithm exhibits particularly low constants. For simplicity, we omit summation terms, as these are negligible and/or dominated by adjacent dot-like operations, and random generation components, as these are fast in theory and practice and the total quantity of random entries generated is
$O\left(\sum_{i=1}^{d}nn_{i}+(m+l)n\mu_{1}+(m+l)n_{d}\right).$
Mat-Mul-Check:
Costs exactly $\mathfrak{m}(b,c,\lambda^{\prime})+\mathfrak{m}(a,b,\lambda^{\prime})+\mathfrak{m}(a,c,\lambda^{\prime})$ (plus vector equality checks, which are negligible in comparison).
Get-Checked-Cum-Prods:
By the above, costs
$\sum_{i=1}^{k-1}\mathfrak{m}(a_{i+1},a_{i+2},\lambda^{\prime})+\mathfrak{m}(a_{i},a_{i+1},\lambda^{\prime})+\mathfrak{m}(a_{i},a_{i+2},\lambda^{\prime})$
for the client and $\sum_{i=1}^{k}\mathfrak{m}(a_{1},a_{i+1},a_{i+2})$ for the server. Note that the first index in both sums is dominant in all uses.
Right-Hidden-Init:
@@
By the above, costs $\sum_{i=0}^{d-1}\mathfrak{m}(s,n_{i},n_{i+1})$ for the server and $\mathfrak{m}(n_{0},n_{1})+\mathfrak{m}(s,n_{0})+\mathfrak{m}(s,n_{1})+\sum_{i=0}^{k-2}\mathfrak{m}(n_{i+1},n_{i+2},\lambda^{\prime})+\mathfrak{m}(n_{i},n_{i+1},\lambda^{\prime})+\mathfrak{m}(n_{i},n_{i+2},\lambda^{\prime})$ for the client.
Right-Hidden-Mult:
Computation of $V^{\prime}$ costs $\mathfrak{m}{n,n_{d},o}+\sum_{i=1}^{d}no\tilde{\mathfrak{d}}(n_{i-1}\mu_{i})$, and computation of $MV^{\prime}$ costs $\mathfrak{m}{s,n_{d},o}+\sum_{i=1}^{d}so\tilde{\mathfrak{d}}(n_{i-1}\mu_{i})$. The product call costs $\mathfrak{m}(s,n,o)$ for the server and $\mathfrak{m}(s,n,\lambda^{\prime})+\mathfrak{m}(n,o,\lambda^{\prime})+\mathfrak{m}(s,o,\lambda^{\prime})$ for the client.
Hidden-Init:
Calculating the cumulative products costs $\sum_{i=0}^{k-2}\mathfrak{m}(n_{i+1},n_{i+2},\lambda^{\prime})+\mathfrak{m}(n_{i},n_{i+1},\lambda^{\prime})+\mathfrak{m}(n_{i},n_{i+2},\lambda^{\prime})$ for the client and $\sum_{i=1}^{k}\mathfrak{m}(n_{1},n_{i+1},n_{i+2})$ for the server. The right-hidden steps are invoked with $s=\sum_{i=1}^{d}n_{i},o=m$. Computation of $A^{\prime}$ costs $\mathfrak{m}(m,n_{d},n)+\sum_{i=0}^{d-1}mn\tilde{\mathfrak{d}}(\mu_{i}n_{i-1})$ for the client.
Hidden-Mult:
Joint computation of $B^{\prime}$ and $AB^{\prime}$ costs $\mathfrak{m}(l,n_{d},n+m)+\sum_{i=0}^{d-1}l(n+m)\tilde{\mathfrak{d}}(\mu_{i}n_{i-1})$. The server pays $\mathfrak{m}(\sum_{i=1}^{d}n_{i},n,l)$ for the in-the-clear computation of the $L$-$(B+B^{\prime})$ products. The client then pays $\mathfrak{m}(m,n_{d},l)+\sum_{i=0}^{d-1}ml\tilde{\mathfrak{d}}(\mu_{i}n_{i-1})$. In the final line, the server pays $\mathfrak{m}(m,n,l)$ and the client pays $2\mathfrak{a}(ml)$.
@@ Appendix 0.B Benchmarks for power-of-two $n$
$n$ Local (s) Client Pre. (s) Server Pre. (s) Client (s) Server (s) Client R. Total R.
@@
Figure 3: Variant of Table 1 without offset $n$. Note that some server time (operand transfer between client in server) is attributed to the client, hence server times slightly below naive time for small operand sizes.
@@
Figure 4: Variant of Table 2 without offset $n$.

### v3

@@ 0.A.1.1 Matrix Multiplication
One natural case is $m=n=l$ and one online step (i.e. a once-off matrix-matrix multiplication). Here, total client cost is $O\left(n^{2}\left(n^{\epsilon}+\nu^{\omega-2}\right)\right)$ and total server cost is, apart from one $n\times n$-by-$n\times n$ matrix multiplication, $O(n^{\omega}\delta^{\omega-2})$. In particular, our protocol is $\left(n^{\epsilon}+\nu^{\omega-2},1+O(\delta^{\omega-2})\right)$-efficient, resulting in a multiplicative $n^{1-\epsilon}$ saving for the client (ignoring the additive term in security parameter) and small overhead for the server.
@@ 0.A.1.2 Online Matrix-Vectors
Another natural case is $m=n,l=1$, and $k$ online steps. Here, setup cost is the same $O\left(n^{2}\left(n^{\epsilon}+\nu^{\omega-2}\right)\right)$ for the client and $O\left(\delta n^{\omega}\right)$ for the server, which amortize across $n^{\epsilon}$ and $\delta n$ online rounds, respectively. The per-multiplication online cost is $O(n(n^{\epsilon}+\nu))$ for the client and, apart from one $n\times n$ matrix-vector multiplication, $O(\delta n^{2})$ for the server. We therefore obtain (after amortization) the same $\left(n^{\epsilon}+\nu,1+O(\delta^{\omega-2})\right)$-efficiency guarantee and small overheads. Note that the amortization occurs in a sublinear number of online steps; this holds because the initial multiplication benefits from fast matrix multiplication, while the online steps do not.
@@ 0.A.2 Per-Operation Analysis
We analyze the performance componentwise and with precise constants, as our protocol exhibits particularly low constants. For simplicity, we omit steps with negligible performance impact.
@@
Let $\tilde{\mathfrak{d}}(a,x)$ be the cost of multiplying a $a\times?$ matrix by a length-??, $x$-sparse vector (considered to be independent of the ??) and $\mathfrak{m}(a,b,c)$ be the cost of multiplying two matrices, one of which is $a\times b$ and the other of which is $b\times c$. All operations are over our working ring $R$.
Protocol 7:
Costs $\sum_{i=2}^{d}\mathfrak{m}(a_{1},a_{i},\lambda^{\prime})+\mathfrak{m}(a_{i},a_{i+1},\lambda^{\prime})+\mathfrak{m}(a_{1},a_{i+1},\lambda^{\prime})$, plus equality checks. Note that the third multiplication can be blocked with either of the two previous.
Protocol 7, Initialization line 2:
One way to order these calculations is to, for every $i\in[d]$, calculate
$L_{j}^{\top}L_{j-1}^{\top}\dots L_{1}^{\top}L_{1}L_{2}\dots L_{i}=\langle L_{j}^{\top}L_{j-1}^{\top}\dots L_{1}^{\top}L_{1}L_{2}\dots L_{i-1}\rangle L_{i},$
for every $j\in[i-1]$ and then calculate
$L_{i}^{\top}L_{j-1}^{\top}\dots L_{1}^{\top}L_{1}L_{2}\dots L_{i}=\langle L_{i-1}^{\top}L_{j-1}^{\top}\dots L_{1}^{\top}L_{1}L_{2}\dots L_{i}\rangle^{\top}L_{i}.$
The calculations in the inner loop can be batched. This gives cost
$\sum_{i=1}^{d}\left[\mathfrak{m}\left(\sum_{j=1}^{i-1}n_{j},n_{i-1},n_{i}\right)+\mathfrak{m}\left(n_{i},n_{i-1},n_{i}\right)\right]$
for the server, and
$\sum_{i=1}^{d}$ $\left[\mathfrak{m}\left(\sum_{j=1}^{i-1}n_{j},n_{i-1}+n_{i},\lambda^{\prime}\right)+2\mathfrak{m}\left(n_{i-1},n_{i},\lambda^{\prime}\right)+\right]+\mathfrak{m}\left(n_{i},n_{i}+n_{i-1},\lambda^{\prime}\right)$
for the client checks.
Protocol 7, Initialization line 4:
This costs $\mathfrak{m}(m,n_{d},n)+m\sum_{i=0}^{d-1}\tilde{\mathfrak{d}}(n,n_{i}\mu_{i+1})$.
Protocol 7, Initialization line 6:
@@
This costs the server $\sum_{i=1}^{d}\mathfrak{m}(m,n_{i-1},n_{i})$, and checking costs the client $\sum_{i=1}^{d}\mathfrak{m}(m,n_{i-1}+n_{i},\lambda^{\prime})+\mathfrak{m}(n_{i-1},n_{i},\lambda^{\prime})$.
Protocol 7, Initialization line 7:
This costs
$\sum_{i=1}^{d}\mathfrak{m}(m,n_{d},n_{i})+\sum_{j=0}^{d-1}m\tilde{\mathfrak{d}}(n_{i},\mu_{j+1}n_{j})$
Protocol 7, Multiplication line 2:
This costs $\mathfrak{m}(l,n_{d},n)+l\sum_{i=0}^{d-1}\tilde{\mathfrak{d}}(n,n_{i}\mu_{i+1})$.
Protocol 7, Multiplication line 3:
This costs $\mathfrak{m}(l,n_{d},m)+l\sum_{i=0}^{d-1}\tilde{\mathfrak{d}}(m,n_{i}\mu_{i+1})$.
Protocol 7, Multiplication line 4:
This costs the server $\mathfrak{m}(m,n,l)$ for the first product (note that this is exactly the naive multiplication cost, and is the only term of this order), and $\sum_{i=1}^{d}\mathfrak{m}(l,n_{i-1},n_{i})$ for the rest by computing partial products.
Protocol 7, Multiplication line 5:
This costs $\mathfrak{m}(m,n_{d},l)+m\sum_{i=0}^{d-1}\tilde{\mathfrak{d}}(l,n_{i}\mu_{i+1})$.
@@ Appendix 0.B Benchmarks for power-of-two $n$
$n$ Local (s) Client Init. (s) Server Init. (s) Client (s) Server (s) Client R. Total R.
@@
Table 3: Variant of Table 1 without offset $n$. Note that some server time is conservatively attributed to the client in simulation, so server times are slightly below naive time for small operand sizes.
@@
Table 4: Variant of Table 2 without offset $n$.